OpenStudy (anonymous):

sam spent $54.60 on games for her little brother. she spend 25% more on games for herself. what was the total cost brought for herself and her little brother?

OpenStudy (anonymous):

Ok so first you have to find what 25% of 54.60 is: we will make that stand for x x/54.60=25/100 cross multiply 54.60(25)=1365 100(x)=100x 1365=100x divide both sides by x x=13.65 so you know that 25% of 54.60 is $13.65 Because she is spending more you add this to the 54.60 54.60+13.65=$68.25

OpenStudy (anonymous):

Make sense?